OpenStudy (anonymous):

i need help please!!!!!!!!!!! find an equation of the line having the given slope and containing the given points slope:-6; through:(6,-7)

pooja195 (pooja195):

y-y1=m(x-x1) m=slope (x1,y1)

OpenStudy (anonymous):

i just need help knowing how to set it up

pooja195 (pooja195):

ok m=slope (in the eqaution) so just plug in the number given for slope into the place of m

pooja195 (pooja195):

Get it?

OpenStudy (anonymous):

let me try to set t up myself

pooja195 (pooja195):

ok :)

OpenStudy (anonymous):

y+7=-6x+36?

pooja195 (pooja195):

Thats right

OpenStudy (anonymous):

okay

pooja195 (pooja195):

Do you get it though ?

OpenStudy (anonymous):

y=-6x+29

pooja195 (pooja195):

Good :) y+7=-6(x-6) y+7=-6x+36 y=-6x+29 yep thats correct good work :)

OpenStudy (anonymous):

cool, thanks

OpenStudy (anonymous):

what if the slope is a fraction?

pooja195 (pooja195):

Still the same :) say your problem was slope =1/7 (6,5) y-5=1/7(x-6)

pooja195 (pooja195):

The way you plug things in will remain the same

OpenStudy (anonymous):

okay thanks so much

pooja195 (pooja195):

:)